OpenStudy (anonymous):

(a^1/2b)^1/2(ab^1/2) This is hard..

OpenStudy (anonymous):

\[\huge \frac{\frac{a}{b}}{\frac{c}{d}} \implies \frac{a \cdot d}{b \cdot c}\]

OpenStudy (mathmath333):